# Parameter estimation of gravitational-wave signals with frequency-dependent antenna responses and higher-modes

### v1

@@ 2.2 Sampling the Likelihood
$p(\Theta|d):=\frac{\mathcal{L}(\Theta|d)p(\Theta)}{\mathcal{Z}}$ (11)
[…]
1 $>$ likelihood = bilby.gw.likelihood.GravitationalWaveTransientNextGeneration(interferometers=ifos, waveform_generator=waveform_generator, priors=priors_basic, distance_marginalization=False, phase_marginalization=False, time_reference=‘geocent’, earth_rotation_time_delay = True, earth_rotation_beam_patterns = True, finite_size = True,)’
We consider a frequency band of 5 Hz to 2048 Hz and so we set the sampling rate to 4096 Hz. For a GW170817-like signal, the dominant mode of radiation $h_{(+,\times)}^{2,2}$, lasts in-band for approximately 2 hours, requiring waveform evaluations at $10^{7}$ points. We need to calculate the waveform for each azimuthal mode m which takes around 10s on a single CPU. Computing the waveform for three azimuthal modes (m=2,3,4) takes approximately 30 seconds on a single CPU. Sampling the likelihood using a nested sampler like dynesty requires around $10^{8}$ waveform evaluations for high SNRs and is computationally prohibitive.
To accelerate likelihood evaluations, we employ two approximations: multibanding and relative binning, described in subsequent sections.
@@ 2.2.1 The Multibanding Approximation
The multibanding approximation [30] is a form of adaptive sampling that divides the entire frequency range into overlapping frequency bands where the start and end frequencies depend on the sequence of durations controlled by an accuracy factor. The duration should be computed using the highest azimuthal mode. This approximation was used to do parameter estimation of signals in Cosmic Explorer containing the dominant $h_{(+,\times)}^{2,2}$ mode [21]. However unlike [21], we compute the antenna response at every banded frequency point instead of every 4 seconds, making the algorithm more accurate without trading-off performance. We implement this approximation as MBGravitationalWaveTransientNextGeneration, a subclass of MBGravitationalWaveTransient in bilby.gw.likelihood.
1$>$ search_waveform_generator_MB = bilby.gw.WaveformGenerator(duration=duration, sampling_frequency=sampling_frequency, frequency_domain_source_model=bilby.gw.source.binary_black_hole_individual_modes_frequency_sequence, parameter_conversion=bilby.gw.conversion.convert_to_lal_binary_black_hole_parameters, waveform_arguments=dict(waveform_approximant=waveformname, reference_frequency=reference_frequency, minimum_frequency=minimum_frequency, mode_array=mode_array))
2$>$ likelihood_MB = bilby.gw.likelihood.MBGravitationalWaveTransientNextGeneration(interferometers=ifos, waveform_generator=search_waveform_generator_MB, priors=priors_MB, reference_chirp_mass=detector_frame_chirp_mass-1e-5, linear_interpolation=True, distance_marginalization=False,phase_marginalization=False, time_reference=‘geocent’, accuracy_factor = 5, highest_mode = 4, earth_rotation_time_delay = True, earth_rotation_beam_patterns = True, finite_size = True)’
The source model binary_black_hole_individual_modes_frequency_sequence can generate waveform polarizations for every azimuthal mode given an arbitrary frequency array. The reference_chirp_mass is an approximate value of chirp mass required for multibanding. The highest_mode should be set to the highest azimuthal mode. The accuracy_factor is a tunable parameter which controls the number of bands and hence the accuracy of the approximation (Referred to $L$ in [30]). A value of 5 results in errors in log-likelihood less than unity. This approximation due to adaptive frequency binning needs only $10^{5}$ evaluations per waveform resulting in a speed-up of $\mathcal{O}$($10^{2}$).
@@ 2.2.2 The Relative Binning Approximation
The relative binning technique accelerates parameter estimation by leveraging the smooth variation of waveform ratios across neighboring points in the parameter space. The waveform corresponding to the maximum likelihood estimate is selected as the fiducial waveform, serving as the reference for constructing approximations. Waveforms at nearby points in the parameter space are computed by applying piecewise linear modifications to the fiducial waveform, significantly reducing the computational overhead of waveform generation. This approach achieves speed-ups of up to $\mathcal{O}(10^{3})$ with minimal loss of accuracy [39, 40]. This method is implemented as RelativeBinningGravitationalWaveTransientNextGeneration in bilby.gw.likelihood.
1$>$ search_waveform_generator_RB = bilby.gw.WaveformGenerator(duration=duration, sampling_frequency=sampling_frequency, frequency_domain_source_model=bilby.gw.source.lal_binary_neutron_star_relative_binning, parameter_conversion=bilby.gw.conversion.convert_to_lal_binary_neutron_star_parameters, waveform_arguments=dict( waveform_approximant=waveformname, reference_frequency=reference_frequency, minimum_frequency=minimum_frequency))
2$>$ likelihood_RB = bilby.gw.likelihood.RelativeBinningGravitationalWaveTransientNextGeneration(interferometers=ifos, waveform_generator=search_waveform_generator_RB, priors=priors_RB, distance_marginalization=False,phase_marginalization=False, time_reference=‘geocent’, chi=10, epsilon=0.1, fiducial_parameters=injection_parameters, earth_rotation_time_delay=True,earth_rotation_beam_patterns=True, finite_size=True,)’
@@ 2.2.3 Mode-by-mode Relative Binning
For systems that include significant contributions from higher-order modes (like mass-asymmetric systems) or precessional effects, the oscillatory nature of the frequency-domain waveform can challenge the accuracy of standard piecewise-linear approximations. While [40] demonstrates that the relative binning technique remains effective even for asymmetric systems, a more robust method—mode-by-mode relative binning—was introduced by [41] to address the complexities arising from the presence of higher modes. This approach decomposes the waveform into its spherical harmonic modes, $(l,m)$, and applies the relative binning approximation to each mode individually. The full waveform is then reconstructed by summing the approximated contributions from all modes. This refinement improves the accuracy of waveform modeling for systems with significant higher mode contributions or precessional dynamics, while retaining the computational advantages of the relative binning framework [42]. This method has been implemented within bilby.gw.likelihood as RelativeBinningGravitationalWaveTransientNextGenerationModebyMode.
1$>$ search_waveform_generator_RB = bilby.gw.WaveformGenerator(duration=duration, sampling_frequency=sampling_frequency, frequency_domain_source_model=bilby.gw.source.lal_binary_black_hole_relative_binning_individual_modes, parameter_conversion=bilby.gw.conversion.convert_to_lal_binary_neutron_star_parameters, waveform_arguments=dict(waveform_approximant=waveformname, reference_frequency=reference_frequency, minimum_frequency=minimum_frequency, mode_array=mode_array))
2$>$ likelihood_RB = bilby.gw.likelihood.RelativeBinningGravitationalWaveTransientNextGenerationModebyMode(interferometers=ifos, waveform_generator=search_waveform_generator_RB, priors=priors_RB, distance_marginalization=False, phase_marginalization=False, time_reference=‘geocent’, chi=10, epsilon=0.1, fiducial_parameters=injection_parameters, earth_rotation_time_delay = True, earth_rotation_beam_patterns = True, inite_size = True,)’
The source model binary_black_hole_relative_binning_individual_modes can generate waveform polarizations required for relative binning for every azimuthal mode. The accuracy is controlled by the usual parameters chi and epsilon which we set to 10 and 0.1 for our work.
@@ 3 Simulations
For illustrative purposes, we consider two high SNR GW170817-like signals which we shall refer to as GW1 and GW2. The only difference between $\Theta_{\rm GW1}$ and $\Theta_{\rm GW2}$ is the inclination angle. The injected parameters and the priors used for sampling are in Table 1. GW2 has an inclination of $\pi/2$ (commonly referred to as an edge-on configuration) and accumulates more SNR for higher azimuthal mode as seen in Table 2.
Parameter $\Theta_{GW1}$ $\Theta_{GW2}$ Priors
Chirp Mass ($\mathcal{M}_{z}$) 1.20994 $M_{\odot}$ 1.20994 $M_{\odot}$ Uniform(1.20993, 1.20995) $M_{\odot}$
Mass Ratio ($q$) 0.918 0.918 Uniform(0.2, 1)
$\chi_{1}^{z}$ 0 0 Uniform(-0.05, 0.05)
$\chi_{2}^{z}$ 0 0 Uniform(-0.05, 0.05)
Right Asc. (RA) 3.44616 3.44616 Uniform(0, $2\pi$)
Declination (Dec) -0.408084 -0.408084 Cosine($-\pi/2$, $\pi/2$)
Incl. Angle ($\theta_{\textrm{JN}}$) 0.53 $\pi/2$ Sine(0, $\pi/2$)
Pol. Angle ($\psi$) 2.212 2.212 Uniform(0, $\pi$)
Phase ($\phi_{c}$) 5.180 5.180 Uniform(0, $2\pi$)
Time at CE ($t_{CE}$) 1187008882.45 s 1187008882.45 s $t_{CE}$ + Uniform (-0.01, 0.01)s
Lum. Distance ($D_{L}$) 46.395 Mpc 46.395 Mpc Uniform in Vol.(10, 2000) Mpc
@@
We use IMRPhenomXPHM [44] to simulate the GW waveforms containing (2, 2), (3, 2), (3, 3) and (4, 4) which are injected into a "zero-noise" realization of Gaussian noise [45] at the detectors. IMRPhenomXPHM is a black hole waveform class, so the tidal deformability parameters are zero by default. However, we do not expect this to affect our results significantly. We also use aligned spin priors for our analyses. For each detector network we perform 3 inferences: including only the $m=2$ mode; excluding the (4, 4) mode alone, and including all modes. The optimal SNR due to each mode is in Table 2.
@@
(2,2) 1847.92 929.32 76.90 993.54 530.48 28.72
(3,2) 1.56 0.86 0.13 2.94 1.76 0.16
(3,3) 5.92 3.14 0.35 6.31 3.54 0.26
(4,4) 2.70 1.51 0.21 5.67 3.38 0.30
Table 2: Optimal SNR collected by each mode of radiation for GW1 and GW2 in the band 5 Hz - 2048 Hz for a 40 km CE, a 20 km CE and LIGO-India.
@@ 4 Results
We use RelativeBinningGravitationalWaveTransientNextGenerationModebyMode to approximate the likelihood for its computational efficiency. For sampling the 11-dimensional parameter space  [^1] For $m=2$ mode only run, we marginalize love the phase analytically and so it is a 10 parameter run in that case., we utilize dynesty, as implemented in Bilby. This dynamic nested sampling algorithm is particularly well-suited for Bayesian inference in complex parameter spaces. The sampling process generates new samples using the acceptance-walk method. In this approach, all Markov Chain Monte Carlo (MCMC) chains run for the same duration at each iteration, with the chain length dynamically adjusted to maintain a target acceptance rate. The number of accepted steps in each chain is controlled to follow a Poisson distribution with a mean of 20. We initialize the sampling with 500 live points and parallelize the process across 16 threads (npool=16). The termination criterion for sampling is set such that the relative change in evidence remains below 0.01.
@@ 4.1 Validation
Figure 1: The histogram of errors in log-likelihood ratio obtained using 2500 samples drawn from the relative binning likelihood in a 40 km CE. For multibanding we use an accuracy_factor of 1 and enabled linear interpolation. For relative binning we set chi to 10 and epsilon to 0.1.
We evaluate the accuracy of the methods described in the previous section by selecting 2,500 downsampled points from the posterior distribution of GW1, observed in a 40 km CE. For each point, the exact log-likelihood ratio is calculated using GravitationalWaveTransientNextGeneration. As CE is the most sensitive detector considered in this study, the differences in log-likelihood ratios are expected to be even smaller for other detectors. The chosen signal has a high signal-to-noise ratio (SNR) and includes higher-order multipole moments, making it well-suited for this analysis. Additionally, we also validate the multibanding approximation. In both cases, the difference in log-likelihood ratios is found to be less than unity [46], as illustrated in Figure 1, confirming the robustness of our results. The errors in the log-likelihood ratio decrease as the signal-to-noise ratio (SNR) decreases, as expected. This reduction in likelihood errors is less pronounced for multibanding compared to relative binning, as the latter does not strictly account for frequency-dependent antenna response effects. The current scheme to determine frequency bins in relative binning is solely based on the Post-Newtonian inspiral phase, but it could be easily extended to incorporate the effects of Earth’s rotation and finite detector sizes, because their oscillatory scales in frequency domain can be easily estimated and frequency bins can be constructed so that those effects are accurately captured by the interpolation. The values of chi and epsilon are chosen such that the log-likelihood errors for the highest SNR signal remain below unity.
@@ 4.2 Inference using 1 Cosmic Explorer
Figure 2: Inferred posteriors of intrinsic parameters of GW1 and GW2 in a 40km CE. The green posteriors use all modes for analysis, the orange posterior ignore modes with m = 3 and the blue ones ignore m = 3 and m = 4 modes.
@@
Figure 3: Inferred posteriors of extrinsic parameters of GW1 and GW2 in one 40km CE. The green posteriors use all modes for analysis, the orange posterior ignore modes with m = 3 and the blue ones ignores m = 3 and m = 4 modes.
@@ 4.2.1 Intrinsic Parameters
The inclusion of higher radiation modes for GW1 has a minimal impact on the intrinsic parameters due to the low signal-to-noise ratios (SNRs) associated with these modes (see Table 2). In the case of a perfectly face-on orientation ($\theta_{\rm JN}=0$), the spin-weighted spherical harmonics vanish for modes where m $\neq$ 2. Given that GW1 has a very small inclination, the effects of higher-order modes are significantly suppressed compared to GW2, as reflected in the posterior distributions.
@@ 4.2.2 Extrinsic Parameters
For the extrinsic parameters as well, the inclusion of higher modes has a greater impact on GW2 than on GW1. However, the posterior distribution of the polarization angle ($\psi$) presents an interesting case. When higher modes are not included in GW1, the recovered posterior follows the prior distribution (see 3(a)). In contrast, for GW2 without higher modes, the posterior exhibits a bimodal distribution (see 3(b)). A GW with its dominant l,m = 2,2 mode for the inspiral portion in a single detector can be written as,
@@
For small $\theta_{\texttt{JN}}$ we can write 14 as,
$D_{\textrm{eff}}=D_{\textrm{L}}(1-\theta_{\textrm{JN}}^{2})^{-1/2}(F_{+}^{2}+F_{\times}^{2})^{-1/2}+\mathcal{O}(\theta_{\textrm{JN}}^{2})$ (14)
@@
So the amplitude does not depend on the polarization angle to linear order in the inclination. This is because when the orbit of the binary is a circle in the sky, that circle is invariant under rotations on the plane of the sky. This is true even beyond the long-wavelength limit. The polarization angle enters equation 12 through the antenna responses and we have shown that the amplitude is independent of $\psi$. The only other place where $\psi$ comes in equation 12 is via the argument of the cosine term. For small $\theta_{\texttt{JN}}$, the term $\tan^{-1}\Big{(}\frac{F_{\times}}{F_{+}}\frac{2\cos\theta_{\textrm{JN}}}{1+\cos^{2}\theta_{\textrm{JN}}}\Big{)}$ reduces to $2\psi+f(\theta,\phi)$. So $\phi_{c}$ and $\psi$ are completely degenerate. Physically this is because the phase of the binary on the circle in the sky is degenerate with an overall rotation of that circle. For the m=2 mode  [^2] Note that we club the modes with the same azimuthal number (m) as they have the same detector response and hence we do not have any results with only the dominant mode of radiation (i.e. l=2, m=2). However we expect results due to only the (2,2) mode and (2,2),(3,2) modes to be identical as the SNR due to the (3,2) mode is negligible 3. only run, it is not possible to recover any information about $\psi$ unless and until we have higher modes or a network sensitive enough to detect quadratic terms in $\theta_{\rm JN}$. For a face-on system as discussed earlier, the spin-weighted spherical harmonics (${}_{-2}Y_{lm}$s) vanish except for m=$\pm$2. So without a m=3, m=4 $\phi_{c}$ and $\psi$ in the waveform’s phase would be degenerate  [^3] We do not show the posteriors of $\phi_{c}$. However we sample $\phi_{c}$ for runs containing m=3 and m=4 modes. For runs containing only the m=2 mode we analytically marginalize over $\phi_{c}$..
For GW2, the same cannot be said as it is completely edge-on. In this limit $\tan^{-1}\Big{(}\frac{F_{\times}}{F_{+}}\frac{2\cos\theta_{\textrm{JN}}}{1+\cos^{2}\theta_{\textrm{JN}}}\Big{)}$ vanishes and the argument of the cosine term in equation 12 is independent of the geometric parameters and there exists no degeneracy with $\phi_{c}$. The amplitude depends only on $|F_{+}|$ which is invariant under the transformation $\psi\hookrightarrow\psi+\frac{\pi}{2}$ which is seen in 3(b). This is infact a more general degeneracy that exists in the polarization angle [47] when higher azimuthal modes are ignored.

### v2

@@ 2.2 Sampling the Posterior
$p(\Theta|d):=\frac{\mathcal{L}(\Theta)p(\Theta)}{\mathcal{Z}}$ (13)
The prior denoted by $p(\Theta)$ is predetermined, and the evidence $\mathcal{Z}$, the integral of the numerator over all parameters ($\Theta$), is a normalization constant. The likelihood ($\mathcal{L}(\Theta)$) assuming that noise is stationary and gaussian is given by $\mathcal{L}(\Theta):=p(d|\Theta)\propto\exp(-\frac{1}{2}{\rm\big<}d-h|d-h{\rm\big>})\propto\exp({\rm\big<}h|d{\rm\big>}-{\rm\big<}h|h{\rm\big>/2})$ where ${\rm\big<}a|b{\rm\big>}=4\penalty 10000\ \text{Re}\int_{0}^{\infty}\frac{a^{*}(f)b(f)}{S_{n}(f)}df$ and $S_{n}(f)$ is the one-sided power spectral density.
We consider a frequency band of 5 Hz to 4096 Hz, and so we set the sampling rate to 8192 Hz. For a GW170817-like signal, the dominant mode of radiation $h_{(+,\times)}^{2,2}$ lasts in-band for approximately 2 hours, requiring waveform evaluations at $10^{7}$ points. We need to calculate the waveform for each azimuthal mode m, which takes around 10s (assuming IMRPhenomXPHM waveform model) on a single CPU. Computing the waveform for three azimuthal modes (m=2,3,4) takes approximately 30 seconds on a single CPU. Sampling the likelihood using a nested sampler like dynesty requires around $10^{8}$ waveform evaluations for high SNRs and is computationally prohibitive. To accelerate likelihood evaluations, we employ two approximations: multibanding and relative binning, described in subsequent sections.
@@ 2.2.1 The Multibanding Approximation
The multibanding approximation [41, 46, 43, 30] is a form of adaptive sampling that divides the entire frequency range into overlapping frequency bands where the start and end frequencies depend on the sequence of durations controlled by an accuracy factor. The duration should be computed using the highest azimuthal mode because, at a fixed frequency, the time to merger for a higher azimuthal mode is longer. This approximation was used to do parameter estimation of signals in Cosmic Explorer containing the dominant $h_{(+,\times)}^{2,2}$ mode [21]. However, unlike [21], we compute the antenna response at every banded frequency point instead of every 4 seconds, making the algorithm more accurate without trading off performance.
We develop a new source model that can generate waveform polarizations for every azimuthal mode given an arbitrary frequency array which is necessary for our multibanding implementation. The highest azimuthal mode and a reference chirp mass value is used to calculate the banding times. A tunable parameter controls the number of bands and hence the accuracy of the approximation (Referred to $L$ in [30]). A value of 5 results in log-likelihood errors less than unity (see figure 1). This approximation, due to adaptive frequency binning, needs only $10^{5}$ evaluations per waveform, resulting in a speed-up of $\mathcal{O}$($10^{2}$).
@@ 2.2.2 The Relative Binning Approximation
The relative binning approximation accelerates parameter estimation by leveraging the smooth variation of waveform ratios as a function of frequency across neighboring points in the parameter space in a typical set of posterior distribution. The waveform corresponding to the maximum likelihood estimate is selected as the fiducial waveform, serving as the reference for constructing approximations. Waveforms at nearby points in the parameter space are computed by applying piecewise linear modifications to the fiducial waveform, significantly reducing the computational overhead of waveform generation.
@@ 2.2.3 Mode-by-mode Relative Binning
For systems that include significant contributions from higher-order modes (like mass-asymmetric systems) or precessional effects, the oscillatory nature of the frequency-domain waveform can challenge the accuracy of standard piecewise-linear approximations. While [40] demonstrates that the relative binning technique remains effective even for asymmetric systems, a more robust method—mode-by-mode relative binning—was introduced by [41], [46] to address the complexities arising from the presence of higher modes. This approach decomposes the waveform into its spherical harmonic modes, $(l,m)$, and applies the relative binning approximation to each mode individually. The full waveform is then reconstructed by summing the approximated contributions from all modes. This refinement improves the accuracy of waveform modeling for systems with significant higher mode contributions or precessional dynamics, while retaining the computational advantages of the relative binning framework [42]. This approach achieves required accuracy with $10^{4}$ evaluations per waveform, resulting in speed-ups of up to $\mathcal{O}(10^{3})$ [39, 40].
@@ 3 Simulations
For illustrative purposes, we consider GW170817-like signals at an optimal SNR of 1000 which we shall refer to as GW1 and GW2. The difference between $\Theta_{\rm GW1}$ and $\Theta_{\rm GW2}$ is the inclination angle. GW2 has an inclination of $\pi/2$ (commonly referred to as an edge-on configuration) and accumulates more SNR for higher azimuthal mode as seen in Table 2. The luminosity distance has been scaled so that the optimal SNR is 1000. The duration of the signal is 107 minutes, assuming the signal enters the CE band at 5Hz. The injected parameters and the priors used for sampling are in Table 1.
Parameter Injected Value Priors
Chirp Mass ($\mathcal{M}_{z}$) 1.22052 (GW1); 1.20983 (GW2) $M_{\odot}$ $\mathcal{M}^{\textrm{inj}}_{z}$ + Uniform($10^{-5},10^{-5}$) $M_{\odot}$
Mass Ratio ($q$) 0.918 Uniform(0.2, 1)
$\chi_{1}^{z}$ 0 Uniform(-0.05, 0.05)
$\chi_{2}^{z}$ 0 Uniform(-0.05, 0.05)
Right Asc. (RA) 3.44616 Uniform(0, $2\pi$)
Declination (Dec) -0.408084 Cosine($\pi/2$, $\pi/2$)
Incl. Angle ($\theta_{\textrm{JN}}$) 0.53 (GW1); $\pi/2$ (GW2) Sine(0, $\pi/2$)
Pol. Angle ($\psi$) 2.212 Uniform(0, $\pi$)
Phase ($\phi_{c}$) 5.180 Uniform(0, $2\pi$)
Time at CE ($t_{CE}$) 1187008882.45 s $t_{CE}$ + Uniform(-0.003, 0.003)s
Lum. Distance ($D_{L}$) 86.35 Mpc (GW1); 46.95 Mpc (GW2) Uniform in Vol.(10, 100) Mpc
@@
We use IMRPhenomXPHM [44] to simulate the GW waveforms containing (2, 2), (3, 2), (3, 3) and (4, 4) which are injected into a "zero-noise" realization of Gaussian noise [45] at the detectors. IMRPhenomXPHM is a black hole waveform class, so the tidal deformability parameters are zero by default. We also use aligned spin priors for our analyses. For each detector network we perform 3 inferences: including only the $m=2$ mode; excluding the (4, 4) mode alone, and including all modes. The optimal SNR due to each mode is in Table 2.
@@
(2,2) 1000 503 41.6 1003 536 28.9
(3,2) 0.85 0.46 0.07 2.96 1.78 0.16
(3,3) 3.21 1.70 0.18 6.37 3.57 0.26
(4,4) 1.47 0.82 0.11 5.72 3.42 0.31
Table 2: Optimal SNR collected by each mode of radiation for GW1 and GW2 in the band 5 Hz - 4096 Hz for a 40 km CE, a 20 km CE and LIGO-India.
@@ 4 Results
We use mode-by-mode relative binning to approximate the likelihood for its computational efficiency. For sampling the 11-dimensional parameter space  [^2] For $m=2$ mode only run, we marginalize the phase analytically and so it is a 10 parameter run in that case., we utilize dynesty, as implemented in Bilby. This dynamic nested sampling algorithm is particularly well-suited for Bayesian inference in complex parameter spaces. The sampling process generates new samples using the acceptance-walk method. In this approach, all Markov Chain Monte Carlo (MCMC) chains run for the same duration at each iteration, with the chain length dynamically adjusted to maintain a target acceptance rate. The termination criterion for sampling is set such that the relative change in evidence remains below 0.01.
@@ 4.1 Validation
Figure 1: The histogram of errors in log-likelihood ratio obtained using 500 samples drawn from the relative binning posterior of GW1 and GW2 in a 40 km CE. For multibanding, we use an accuracy_factor of 1 and enabled linear interpolation. For relative binning, we set chi to 10 and epsilon to 0.1. It is worth noting that the histogram can shift left and right on the horizontal scale, which depends on the choice of approximation parameter. We only require that the errors are less than unity for unbiased PE.
We evaluate the accuracy of the methods described in the previous section by selecting 500 points from the posterior distributions of GW1 and GW2, observed in a 40 km CE. For each point, the exact log-likelihood ratio is calculated using the exact likelihood. As CE is the most sensitive detector considered in this study, the differences in log-likelihood ratios are expected to be even smaller for other detectors. The chosen signal has a high SNR and includes higher-order modes, making it well-suited for this analysis. Additionally, we also validate the multibanding approximation. We tune the adjustable parameters for the approximate likelihoods such that in both cases, the difference in log-likelihood ratios is found to be less than unity which satisfies the criterion in [46], as illustrated in Figure 1, confirming the robustness of our results. In Figure 1, relative binning appears more accurate than multibanding; however, this is an artifact of the chosen accuracy-controlling parameters. The parameters chosen for multibanding perform worse when higher modes are present. The exact reason for this is left for future investigation. The current scheme to determine frequency bins in relative binning is solely based on the Post-Newtonian inspiral phase, but it can be extended to incorporate the effects of Earth’s rotation and finite detector sizes, because their oscillatory scales in the frequency domain can be easily estimated. Frequency bins can be constructed so that the interpolation accurately captures those effects.
@@ 4.2 Inference using 1 Cosmic Explorer
Figure 2: Inferred posteriors of intrinsic parameters of GW1 and GW2 in a 40km CE. The green posteriors use all modes for analysis, the orange posteriors ignore modes with m = 3 and the blue ones ignore m = 3 and m = 4 modes.
@@
Figure 3: Inferred posteriors of extrinsic parameters of GW1 and GW2 in one 40km CE. The green posteriors use all modes for analysis, the orange posterior ignores modes with m = 3, and the blue ones ignore m = 3 and m = 4 modes.
@@ 4.2.1 Intrinsic Parameters
The inclusion of higher radiation modes for GW1 has a minimal impact on the intrinsic parameters due to the low signal-to-noise ratios (SNRs) associated with these modes (see Table 2). In the case of a perfectly face-on orientation ($\theta_{\rm JN}=0$), the spin-weighted spherical harmonics vanish for modes where m $\neq$ 2. GW1 has an inclination of 0.53, which results in the significant suppression (c.f. table 2) of the effects of higher-order modes. For GW2, ignoring higher-order modes leads to significant biases in the recovered masses as in figure 2(b).
@@ 4.2.2 Extrinsic Parameters
For the extrinsic parameters as well, the inclusion of higher modes has a greater impact on GW2 than on GW1. As seen in the case of intrinsic parameters, exclusion of the m = 3 mode leaders to biased parameter recovery for GW2 (see figure 3(b)).
The posterior distribution of the polarization angle ($\psi$) presents an interesting case. When higher modes are not included in GW1, the recovered posterior follows the prior distribution (see 3(a)). In contrast, for GW2 without higher modes, the posterior exhibits a bimodal distribution (see 3(b)). A GW with its dominant l,m = 2,2 mode for the inspiral portion in a single detector can be written as,
@@
For small $\theta_{\texttt{JN}}$ we can write 16 as,
$D_{\textrm{eff}}=D_{\textrm{L}}(1-\theta_{\textrm{JN}}^{2})^{-1/2}(F_{+}^{2}+F_{\times}^{2})^{-1/2}+\mathcal{O}(\theta_{\textrm{JN}}^{4})$ (16)
@@
[…]
For GW2, the same cannot be said as it is completely edge-on. In this limit $\tan^{-1}\Big(\frac{F_{\times}}{F_{+}}\frac{2\cos\theta_{\textrm{JN}}}{1+\cos^{2}\theta_{\textrm{JN}}}\Big)$ vanishes and the argument of the cosine term in equation 14 is independent of the geometric parameters and there exists no degeneracy with $\phi_{c}$. The amplitude depends only on $|F_{+}|$ which is invariant under the transformation $\psi\hookrightarrow\psi+\frac{\pi}{2}$ which is seen in 3(b). This is a more general degeneracy that exists in the polarization angle [47] when higher azimuthal modes are ignored.
4.2.3 Inference with constant antenna response
We perform parameter estimation while neglecting Earth’s rotation and detector size effects, assuming a static antenna response throughout the signal duration. Specifically, we fix the antenna response to its value at the time of merger. For GW1 (see figure 4(a)), where higher-order modes are negligible, we nonetheless observe significant biases in the recovered parameters. This is consistent with [21].
In contrast, for GW2 (see figure 4(b)), we recover the intrisic parameters accurately. It is important to note that all simulated signals in this work include contributions from all relevant modes. However, as in the case of GW1, we observe noticeable biases in both the inclination angle and luminosity distance.
(a) GW1
(b) GW2
Figure 4: Inferred posteriors of GW1 and GW2 in one 40km CE with effects due to Earth’s rotation and size of the detector turned off.